# TimeDRL: Disentangled Representation Learning for Multivariate Time-Series

### 07 Dec 2023 v1

@@ Abstract
Multivariate time-series data in numerous real-world applications (e.g., healthcare and industry) are informative but challenging due to the lack of labels and high dimensionality. Recent studies in self-supervised learning have shown their potential in learning rich representations without relying on labels, yet they fall short in learning disentangled embeddings and addressing issues of inductive bias (e.g., transformation-invariance). To tackle these challenges, we propose TimeDRL, a generic multivariate time-series representation learning framework with disentangled dual-level embeddings. TimeDRL is characterized by three novel features: (i) disentangled derivation of timestamp-level and instance-level embeddings from patched time-series data using a [CLS] token strategy; (ii) utilization of timestamp-predictive and instance-contrastive tasks for disentangled representation learning, with the former optimizing timestamp-level embeddings with predictive loss, and the latter optimizing instance-level embeddings with contrastive loss; and (iii) avoidance of augmentation methods to eliminate inductive biases, such as transformation-invariance from cropping and masking. Comprehensive experiments on 6 time-series forecasting datasets and 5 time-series classification datasets have shown that TimeDRL consistently surpasses existing representation learning approaches, achieving an average improvement of forecasting by 57.98% in MSE and classification by 1.25% in accuracy. Furthermore, extensive ablation studies confirmed the relative contribution of each component in TimeDRL’s architecture, and semi-supervised learning evaluations demonstrated its effectiveness in real-world scenarios, even with limited labeled data.
@@ I Introduction
Self-supervised learning (SSL) is a prominent method within unsupervised representation learning, which captures generalizable representations from unlabeled data with pretext tasks. This is beneficial in downstream scenarios with limited labeled records from the generic knowledge of pre-trained models, which has been demonstrated in extracting valuable representations in Natural Language Processing (NLP) [7, 8] and Computer Vision (CV) [9, 10, 11]. However, directly applying SSL in the time-series domain suffers from two challenges.
@@
The first challenge in SSL for time-series data is learning disentangled dual-level representations. Existing approaches focus on deriving either timestamp-level [12, 13] or instance-level embeddings [14, 15, 16], but not both at the same time. However, these two types of embeddings serve distinct purposes: timestamp-level embeddings are effective for anomaly detection and forecasting, whereas instance-level embeddings are suited for classification and clustering tasks [17]. Although we can theoretically avoid explicitly deriving instance-level embeddings by extracting them from timestamp-level embeddings using pooling methods (as illustrated in Fig. 1(a)) [12], this approach often results in the anisotropy problem [18, 19, 20], where the embeddings are confined to a narrow cone region in the embedding space, thus limiting their expressiveness. To the best of our knowledge, how to disentangle instance-level embeddings from timestamp-level embeddings in the time-series domain remains an unexplored problem.
@@
To address the above issues, we propose TimeDRL, a generic multivariate Time-series framework with Disentangled Representation Learning for dual-level embeddings. By disentangling timestamp-level and instance-level embeddings (as illustrated in Fig. 1(b)), TimeDRL is applicable across various time-series downstream tasks. We adopt a dedicated [CLS] token to extract the contextualized instance-level embeddings directly from patched timestamp-level data, effectively addressing the anisotropy problem. This method involves aggregating adjacent time steps to form patched time-series data, followed by appending a [CLS] token at the beginning to represent the instance-level representation. Notably, using patched time-series data, as opposed to point-level, allows us to extend the horizon without significantly increasing training costs, ensuring that the [CLS] token captures more comprehensive semantic information. To optimize the encoder network, TimeDRL employs two pretext tasks: a timestamp-predictive task for optimizing timestamp-level embeddings, and an instance-contrastive task for optimizing instance-level embeddings. To mitigate the inductive bias of transformation-invariance, we refrain from directly applying any augmentation to the data in both predictive and contrastive learning tasks, as illustrated in (c) and (d) of Fig. 2. In the timestamp-predictive task, we opt to use reconstruction error on patched time-series data, notably without masking any input data. For the instance-contrastive task, we use dropout layers to introduce variation. By leveraging the inherent randomness of dropout layers, we can generate distinct views of the embeddings from the same input without relying on any external augmentation methods.
In summary, the main contributions of this work are as follows:
@@ II-B Self-Supervised Learning for Time-Series Data
Figure 3: TimeDRL framework. We adopt a Siamese network architecture, using a Transformer Encoder as our encoder $f_{\theta}$, to generate two distinct views of embeddings from the same input. This is achieved by exploiting the inherent randomness of the dropout layers within the encoder, thus eliminating the need for data augmentations. To optimize the encoder, two pretext tasks are employed: the timestamp-predictive task which targets timestamp-level embeddings, and the instance-contrastive task, focusing on instance-level embeddings derived from two different views.
@@ IV-A Disentangled Dual-Level Embeddings
In BERT and RoBERTa, the [CLS] token is employed to capture sentence-level embeddings, which motivated us to adopt it to extract instance-level embeddings in the time-series domain. While it is theoretically possible to derive instance-level embeddings from timestamp-level embeddings through pooling methods (e.g., global average pooling) [12], this approach can lead to the anisotropy problem [18, 19, 20]. This issue arises when the timestamp-level embeddings are restricted to a narrow region in the embedding space, reducing their capacity to effectively capture diverse information. Studies in NLP [27, 40] have indicated that compared to traditional pooling methods, optimizing the [CLS] token through contrastive learning can yield better results. This finding aligns with our experimental outcomes, as detailed in Table VIII.
@@
After receiving the encoder input $\textbf{x}_{enc\_in}$, we pass it into the encoder $f_{\theta}$, which comprises a linear token encoding layer $W_{token}\in\mathbb{R}^{D\times C\cdot P}$, a linear positional encoding layer $W_{pos}\in\mathbb{R}^{D\times(1+T_{p})}$, and a series of self-attention blocks SAs (with $L$ blocks in total). This process yields the final embeddings $\textbf{z}\in\mathbb{R}^{(1+T_{p})\times D}$:
$\textbf{z}=\text{SAs}(W_{token}\textbf{x}_{enc\_in}+W_{pos}).$ (3)
@@ IV-C Instance-Contrastive Task in TimeDRL
In addition, naively using this technique fails to address sampling bias in contrastive learning. Sampling bias occurs when randomly selected negative samples are similar to the positive samples, which is a common scenario in the time-series domain due to the presence of periodic patterns. To mitigate this issue, the proposed method incorporates an additional prediction head along with a stop-gradient operation to prevent model collapse [11, 42]. Furthermore, this negative-free approach in contrastive learning removes the need for large batch sizes, which are usually essential to gather enough negative samples for stable training.
@@ Methods
with $N$ denoting the total number of samples.
@@
TimeDRL
MHCCL CCL SimCLR BYOL TS2Vec TSTCC T-Loss
@@ V-4 Implementation Details
We partition the dataset into three segments: 60% for training, 20% for validation, and 20% for testing, except when a predefined train-test split exists. We employ AdamW [51] optimizer with weight decay. The experiments are conducted on an NVIDIA GeForce RTX 3090 GPU. We use Transformer Encoder as the encoder’s architecture $f_{\theta}$. We design the timestamp-predictive head $p_{\theta}$ using a linear layer, and we design the instance-contrastive head $p_{\theta}$ using a two-layer bottleneck MLP with BatchNorm and ReLU in the middle. In the time-series forecasting task, we incorporate channel-independence alongside patching, a concept also introduced by PatchTST [36]. This method treats multivariate time-series as multiple univariate series, processed collectively by a single model. Although channel-mixing models directly exploit cross-channel data, channel-independence captures cross-channel interactions indirectly through shared weights. We observed that channel-independence significantly enhances performance in time-series forecasting, leading to its integration in our experiments. However, for time-series classification, we found that omitting channel-independence yielded better results.
Figure 4: Semi-supervised learning. To simulate a real-world scenario with limited availability of labeled data, we randomly exclude a portion of the labels in our datasets. In this experiment, ’TimeDRL (FT)’ indicates that we fine-tune the encoder during the downstream tasks.
@@ V-A Linear Evaluation on Time-Series Forecasting
To assess the effectiveness of TimeDRL’s timestamp-level embeddings, we conduct a linear evaluation on time-series forecasting. This involves pre-training the encoder using pretext tasks, then freezing the encoder weights and attaching a linear layer for training on the downstream forecasting task. Following the experimental settings used in SimTS [13], we set various prediction lengths $T\in\{24,48,168,336,720\}$ for datasets like ETTh1, ETTh2, Exchange, and Weather; and $T\in\{24,48,96,228,672\}$ for ETTm1 and ETTm2. The performance of TimeDRL in multivariate forecasting is summarized in Table III, where it shows an average improvement of 57.98% in MSE compared to state-of-the-art methods. Notably, TimeDRL outperforms all baseline methods across all datasets and prediction lengths. The improvement is especially noticeable with longer prediction lengths, indicating TimeDRL’s effectiveness in learning timestamp-level representations for long-term forecasting—a key task in real-world applications. For instance, on the ETTh2 dataset with a prediction length of 720, TimeDRL achieves a 78.96% improvement in MSE. Additionally, to assess the model’s capability with univariate time-series data, we conducted experiments on univariate forecasting. The results, as shown in Table IV, reveal an average improvement of 29.09% in MSE, further validating the versatility of TimeDRL.
@@ V-B Linear Evaluation on Time-Series Classification
To evaluate the efficacy of TimeDRL’s instance-level embeddings, we utilize a linear evaluation approach for time-series classification. Following a similar approach as in our forecasting evaluation, we first train the encoder using self-supervised learning, freeze its weights, and then attach a linear layer for training on classification tasks. The performance of TimeDRL in time-series classification is detailed in Table V, demonstrating an average accuracy improvement of 1.25% over state-of-the-art methods. For datasets where baseline models already achieve around 90% accuracy or higher, TimeDRL consistently maintains comparable results. Notably, on the challenging FingerMovements dataset, where baseline models generally underperform, TimeDRL shows a significant 20.94% improvement in accuracy and a 45.83% enhancement in Cohen’s Kappa coefficient. Furthermore, TimeDRL’s capability with univariate data is evident in the Epilepsy dataset, where it is only marginally lower than the best baseline method by 0.07% in accuracy.
Table VI: Ablation study on pretext tasks. We use the prediction length $T=168$ for ETTh1. The best results are in bold.
Pretext Tasks
ETTh1
FingerMovements
MSE
MAE
ACC MF1 $\kappa$
TimeDRL 0.418 0.427 63.00 63.00 26.06
TimeDRL w/o Timestamp-Predictive
0.423 (-1.2%)
0.430 (-0.58%)
56.00 (-11.11%) 56.00 (-11.11%) 12.04 (-53.8%)
TimeDRL w/o Instance-Contrastive
0.505 (-20.81%)
0.479 (-12.05%)
55.00 (-12.7%) 55.00 (-12.7%) 10.00 (-61.63%)
@@ V-C Semi-supervised learning
The most practical application of self-supervised learning in real-world scenarios is semi-supervised learning, where labeled data are limited, but a vast amount of unlabeled data exists. Traditional supervised learning methods focus solely on limited labeled data, neglecting the untapped potential of unlabeled data. Self-supervised learning shines in this context by enabling the use of extensive unlabeled data for powerful representation learning. We first train an encoder on a large unlabeled dataset to learn rich representations, followed by fine-tuning with limited labeled data alongside the downstream task head. Unlike the linear evaluation in previous sections where the encoder weights are frozen, in this real-world application scenario, the encoder weights are adjusted during fine-tuning. To emulate the situation of limited labeled data availability, we randomly withhold a portion of labels in our datasets. The comparative results of using only labeled data (supervised learning) versus combining both unlabeled and labeled data (TimeDRL with fine-tuning) are illustrated in Fig. 4. The results indicate that incorporating unlabeled data with TimeDRL significantly improves performance, particularly when there is minimal available labeled data. Remarkably, the benefits of TimeDRL’s pretraining stage are evident even when 100% of the labels are available.
@@ V-D1 Pretext Tasks
In TimeDRL, we strategically employ two pretext tasks to optimize both timestamp-level and instance-level embeddings. The timestamp-predictive task focuses on applying loss specifically to timestamp-level embeddings, while the instance-contrastive task targets instance-level embeddings. Our experiment investigates the impact of each task on the effectiveness of representation learning. According to the results in Table VI, it is evident that combining both pretext tasks yields the best performance in both forecasting and classification tasks, highlighting the value of each task in enhancing dual-level embeddings. Notably, the timestamp-predictive task, despite primarily optimizing timestamp-level embeddings, also significantly improves performance in time-series classification tasks, which depend on instance-level embeddings. In scenarios where the instance-contrastive task is employed for instance-level embeddings, an additional 12.50% gain in accuracy is observed when the timestamp-predictive task is also applied. A similar effect is observed in the time-series forecasting task when employing the instance-contrastive task. In time-series forecasting tasks, the addition of the instance-contrastive task to the timestamp-predictive task yields a 17.22% improvement in MSE. These findings underscore the significance of both pretext tasks in the overall efficacy of TimeDRL across different time-series applications.

### 13 Mar 2024 v2

@@ Abstract
Multivariate time-series data in numerous real-world applications (e.g., healthcare and industry) are informative but challenging due to the lack of labels and high dimensionality. Recent studies in self-supervised learning have shown their potential in learning rich representations without relying on labels, yet they fall short in learning disentangled embeddings and addressing issues of inductive bias (e.g., transformation-invariance). To tackle these challenges, we propose TimeDRL, a generic multivariate time-series representation learning framework with disentangled dual-level embeddings. TimeDRL is characterized by three novel features: (i) disentangled derivation of timestamp-level and instance-level embeddings from patched time-series data using a [CLS] token strategy; (ii) utilization of timestamp-predictive and instance-contrastive tasks for disentangled representation learning, with the former optimizing timestamp-level embeddings with predictive loss, and the latter optimizing instance-level embeddings with contrastive loss; and (iii) avoidance of augmentation methods to eliminate inductive biases, such as transformation-invariance from cropping and masking. Comprehensive experiments on 6 time-series forecasting datasets and 5 time-series classification datasets have shown that TimeDRL consistently surpasses existing representation learning approaches, achieving an average improvement of forecasting by 58.02% in MSE and classification by 1.48% in accuracy. Furthermore, extensive ablation studies confirmed the relative contribution of each component in TimeDRL’s architecture, and semi-supervised learning evaluations demonstrated its effectiveness in real-world scenarios, even with limited labeled data. The code is available at https://github.com/blacksnail789521/TimeDRL.
@@ I Introduction
Self-supervised learning (SSL) is a prominent method within unsupervised representation learning, which captures generalizable representations from unlabeled data with pretext tasks. This is beneficial in downstream scenarios with limited labeled records from the generic knowledge of pre-trained models, which has been demonstrated in extracting valuable representations in Natural Language Processing (NLP) [7, 8] and Computer Vision (CV) [9, 10, 11]. However, applying SSL directly in the time-series domain faces two challenges.
@@
The first challenge in SSL for time-series data is learning disentangled dual-level representations. Existing approaches focus on deriving either timestamp-level [12, 13] or instance-level embeddings [14, 15, 16], but not both at the same time. However, these two types of embeddings serve distinct purposes: timestamp-level embeddings are effective for forecasting and anomaly detection, whereas instance-level embeddings are suited for classification and clustering tasks [17]. Although we can theoretically avoid explicitly deriving instance-level embeddings by extracting them from timestamp-level embeddings using pooling methods (as illustrated in Fig. 1(a)) [12], this approach often results in the anisotropy problem [18, 19, 20], where the embeddings are confined to a narrow cone region in the embedding space, thus limiting their expressiveness. To the best of our knowledge, how to disentangle instance-level embeddings from timestamp-level embeddings in the time-series domain remains an unexplored problem.
@@
To address the above issues, we propose TimeDRL, a generic multivariate Time-series framework with Disentangled Representation Learning for dual-level embeddings. By disentangling timestamp-level and instance-level embeddings (as illustrated in Fig. 1(b)), TimeDRL is applicable across various time-series downstream tasks. This method involves aggregating adjacent time steps to form patched time-series data, followed by appending a [CLS] token at the beginning to represent the instance-level representation. Notably, using patched time-series data, as opposed to point-level, allows us to extend the horizon without significantly increasing training costs, ensuring that the [CLS] token captures more comprehensive semantic information. To optimize the encoder network, TimeDRL employs two pretext tasks: a timestamp-predictive task for optimizing timestamp-level embeddings, and an instance-contrastive task for optimizing instance-level embeddings. Specifically, two pretext tasks are designed in a disentangled manner, each of which is specifically tailored to optimize its respective embedding level, ensuring that the optimization of one embedding type does not interfere with the other. To mitigate the inductive bias of transformation-invariance, we refrain from directly applying any augmentation to the data in both predictive and contrastive learning tasks, as illustrated in (c) and (d) of Fig. 2. In the timestamp-predictive task, we opt to use reconstruction error on patched time-series data, notably without masking any input data. For the instance-contrastive task, we use dropout layers to introduce variation. By leveraging the inherent randomness of dropout layers, we can generate distinct views of the embeddings from the same input without relying on any external augmentation methods.
In summary, the paper’s main contributions are as follows:
@@ II-B Self-Supervised Learning for Time-Series Data
Figure 3: TimeDRL framework. The framework is illustrated in two sections: (a) Pre-training Stage and (b) Fine-tuning Stage. In the pre-training stage (a), we adopt a Siamese network architecture with a Transformer encoder as our encoder $f_{\theta}$ to generate two distinct views of embeddings from the same input. This is achieved by exploiting the inherent randomness of the dropout layers within the encoder, thus eliminating the need for data augmentations. The fine-tuning stage (b) demonstrates the application of these embeddings in downstream tasks, showcasing the adaptability of TimeDRL’s pre-trained embeddings for time-series forecasting and classification.
@@ IV-A Disentangled Dual-Level Embeddings
In BERT and RoBERTa, the [CLS] token is employed to capture sentence-level embeddings, which motivated us to adopt it to extract instance-level embeddings in the time-series domain. While it is theoretically possible to derive instance-level embeddings from timestamp-level embeddings through pooling methods (e.g., global average pooling) [12], this approach can lead to the anisotropy problem [18, 19, 20]. This issue arises when the timestamp-level embeddings are restricted to a narrow region in the embedding space, reducing their capacity to effectively capture diverse information. Studies in NLP [27, 40] have indicated that compared to traditional pooling methods, optimizing the [CLS] token through contrastive learning can yield better results. This finding aligns with our experimental outcomes, as detailed in Table VII.
@@
After receiving the encoder input $\textbf{x}_{enc\_in}$, we pass it into the encoder $f_{\theta}$, which comprises a linear token encoding layer $W_{token}\in\mathbb{R}^{D\times C\cdot P}$, a learnable linear positional encoding layer $PE\in\mathbb{R}^{(1+T_{p})\times D}$, and a series of Transformer blocks TBs (with $L$ blocks in total). This process yields the final embeddings $\textbf{z}\in\mathbb{R}^{(1+T_{p})\times D}$:
$\textbf{z}=\text{TBs}(\textbf{x}_{enc\_in}W_{token}^{\top}+PE).$ (3)
@@ IV-C Instance-Contrastive Task in TimeDRL
In addition, naively using this technique fails to address sampling bias in contrastive learning. Sampling bias occurs when randomly selected negative samples are similar to the positive samples, which is a common scenario in the time-series domain due to the presence of periodic patterns. To that end, we focus exclusively on positive samples and remove the use of negative samples to address the challenge of sampling bias. The proposed method incorporates an additional prediction head and integrates a stop-gradient operation to prevent model collapse [11, 42]. Furthermore, this negative-free approach in contrastive learning removes the need for large batch sizes, which are usually essential to gather enough negative samples for stable training.
@@ Methods
with $N$ denoting the total number of samples. The coefficient ranges from $-1$ (total disagreement) to $1$ (perfect agreement), with 00 indicating no agreement beyond chance. Cohen’s Kappa ($\kappa$) is crucial for evaluating classifiers on imbalanced datasets because it adjusts for chance agreement. This reveals when a classifier’s performance is akin to random chance ($\kappa$ close to 00) or when $\kappa$ is negative, indicating performance worse than random, providing critical insights into handling class imbalance.
@@
TimeDRL MHCCL CCL SimCLR BYOL TS2Vec TSTCC T-Loss
@@ V-4 Implementation Details
We partition the dataset into three segments: 60% for training, 20% for validation, and 20% for testing, except when a predefined train-test split exists. We employ AdamW [51] optimizer with weight decay. The experiments are conducted on an NVIDIA GeForce RTX 3070 GPU. We use Transformer encoder as the encoder’s architecture $f_{\theta}$. We design the timestamp-predictive head $p_{\theta}$ using a linear layer, and we design the instance-contrastive head $p_{\theta}$ using a two-layer bottleneck MLP with BatchNorm and ReLU in the middle. In the time-series forecasting task, we incorporate channel-independence alongside patching, a concept also introduced by PatchTST [36]. This method treats multivariate time-series as multiple univariate series, processed collectively by a single model. Although channel-mixing models directly exploit cross-channel data, channel-independence captures cross-channel interactions indirectly through shared weights. We observed that channel-independence significantly enhances performance in time-series forecasting, leading to its integration in our experiments. However, for time-series classification, we found that omitting channel-independence yielded better results.
@@ V-A Linear Evaluation on Time-Series Forecasting
To assess the effectiveness of TimeDRL’s timestamp-level embeddings, we conduct a linear evaluation on time-series forecasting. This involves pre-training the encoder using pretext tasks, then freezing the encoder weights and attaching a linear layer for training on the downstream forecasting task. Following the experimental settings used in SimTS [13], we set various prediction lengths $T\in\{24,48,168,336,720\}$ for datasets like ETTh1, ETTh2, Exchange, and Weather; and $T\in\{24,48,96,228,672\}$ for ETTm1 and ETTm2. The performance of TimeDRL in multivariate forecasting is summarized in Table III, where it shows an average improvement of 58.02% in MSE compared to state-of-the-art methods.
Notably, TimeDRL surpasses the performance of all baselines, including SimTS, despite its contrastive learning objective being specifically designed for forecasting tasks. Although SimTS is effective at predicting the latent representation of future data from historical data, its performance is limited by not accounting for randomness in the data. In contrast, TimeDRL incorporates randomness, allowing it to capture temporal dynamics more effectively and enhance prediction accuracy. The omission of randomness in SimTS results in reduced generalizability, as it fails to account for the inherent variability and unpredictability in real-world scenarios. This is particularly evident with longer prediction lengths. For example, on the ETTh2 dataset with a prediction length of 720, TimeDRL achieves a 78.96% improvement in MSE over SimTS, underscoring its significant advancement in long-term forecasting accuracy. Additionally, to assess the model’s capability with univariate time-series data, we conducted experiments on univariate forecasting. The results, as shown in Table IV, reveal an average improvement of 29.09% in MSE, further validating the versatility of TimeDRL.
Figure 4: Comparison of training time (in seconds) in the pre-training stage on forecasting datasets.
The execution time of TimeDRL, alongside the top-performing baselines SimTS and TS2Vec, is depicted in Fig. 4, utilizing an NVIDIA GeForce RTX 3070 GPU for evaluation. For a fair comparison across all methods, we set the batch size to 32, the number of epochs to 10, and the sequence length $T$ to 512. SimTS and TS2Vec use fast, convolutional-based encoders, while TimeDRL utilizes a Transformer encoder, known for its superior ability to capture temporal dependencies but with longer execution times. To improve efficiency, TimeDRL integrates a patching mechanism, reducing input sequence length from $L$ to $\lfloor(L-P)/S\rfloor+2$, thus lowering computational and memory demands quadratically. Despite TimeDRL’s longer execution time compared to convolutional-based counterparts, its patching mechanism significantly closes the efficiency gap, enhancing its ability to capture complex temporal dependencies without compromising performance.
Figure 5: Semi-supervised learning. To simulate real-world scenarios with limited availability of labeled data, we experiment with different portions of the labels in our datasets. The top sections (a-c) represent time-series forecasting, while the bottom sections (d-f) correspond to time-series classification. ’TimeDRL (FT)’ indicates that we fine-tune the encoder during the downstream tasks.
@@ V-B Linear Evaluation on Time-Series Classification
To evaluate the efficacy of TimeDRL’s instance-level embeddings, we utilize a linear evaluation approach for time-series classification. Following a similar approach as in our forecasting evaluation, we first train the encoder using self-supervised learning, freeze its weights, and then attach a linear layer for training on classification tasks. The performance of TimeDRL in time-series classification is detailed in Table V, demonstrating an average accuracy improvement of 1.48% over state-of-the-art methods. For datasets where baseline models already achieve around 90% accuracy or higher, TimeDRL consistently maintains comparable results. Notably, on the challenging FingerMovements dataset, where baseline models generally underperform, TimeDRL shows a significant 22.86% improvement in accuracy and a 58.13% enhancement in Cohen’s Kappa coefficient. Furthermore, TimeDRL’s capability with univariate data is evident in the Epilepsy dataset, which is only marginally lower than the best baseline method by 0.07% in accuracy.
@@ V-C Semi-supervised learning
The most practical application of self-supervised learning in real-world scenarios is semi-supervised learning, where labeled data are limited, but a vast amount of unlabeled data exists. Traditional supervised learning methods focus solely on limited labeled data, neglecting the untapped potential of unlabeled data. Self-supervised learning shines in this context by enabling the use of extensive unlabeled data for powerful representation learning. We first train an encoder on a large unlabeled dataset to learn rich representations, followed by fine-tuning with limited labeled data alongside the downstream task head. Unlike the linear evaluation in previous sections where the encoder weights are frozen, in this real-world application scenario, the encoder weights are adjusted during fine-tuning. To emulate the situation of limited labeled data availability, we randomly withhold a portion of labels in our datasets. The comparative results of using only labeled data (supervised learning) versus combining both unlabeled and labeled data (TimeDRL with fine-tuning) are illustrated in Fig. 5. The results indicate that incorporating unlabeled data with TimeDRL significantly boosts performance in both forecasting (with MSE) and classification (with accuracy), particularly as the proportion of available labeled data decreases. This trend highlights TimeDRL’s efficiency in utilizing unlabeled data to boost performance as the availability of labeled data reduces. Remarkably, the benefits of TimeDRL’s pre-training stage are evident even when 100% of the labels are available.
@@ V-D1 Pretext Tasks
In TimeDRL, we strategically employ two pretext tasks to optimize both timestamp-level and instance-level embeddings. The timestamp-predictive task focuses on applying loss specifically to timestamp-level embeddings, while the instance-contrastive task targets instance-level embeddings. Our experiment conducts a sensitivity analysis of the lambda parameter in Equation (19) to assess its impact on the effectiveness of representation learning. According to the results in Fig. 6, it is evident that combining both pretext tasks yields the best performance in both forecasting and classification tasks, highlighting the value of each task in enhancing dual-level embeddings. Notably, the instance-contrastive task, despite primarily optimizing instance-level embeddings, significantly improves performance in time-series forecasting tasks, which depend on timestamp-level embeddings. In scenarios where the instance-contrastive task contributes minimally to the overall loss ($\lambda=0.001$), the MSE experiences a substantial increase compared to scenarios where both losses are leveraged equally ($\lambda=1$). A similar effect is observed in the time-series classification task, where neglecting the timestamp-predictive task in favor of the instance-contrastive task ($\lambda=1000$) leads to a significant decrease in accuracy. These findings underscore the significance of both pretext tasks in the overall efficacy of TimeDRL across different time-series applications.
Figure 6: Sensitivity analysis on $\lambda$. When $\lambda$ is smaller, the focus is more on predictive loss ($\mathcal{L}_{P}$); conversely, a larger $\lambda$ shifts the focus towards contrastive loss ($\mathcal{L}_{C}$).